OpenStudy (bekkah323):

Divide and express your results in scientific notation? 6.4*10^-7/8.0*10^-7

OpenStudy (anonymous):

\[\frac{6.4}{8}=?\]?

OpenStudy (bekkah323):

.8

OpenStudy (anonymous):

ok good, and \(\large \frac{10^{-7}}{10^{-7}}=1\) right?

OpenStudy (bekkah323):

yes

OpenStudy (anonymous):

i mean \[\large \frac{10^{-7}}{10^{-7}}=1\]

OpenStudy (anonymous):

so you get \(.8\), which is not in scientific notation until you write it as \(8\times 10^{-1}\)